# Are Colors Quanta of Light for Human Vision? A Quantum Cognition Study of Visual Perception

### v1

@@ 4 Categorical Perception and Quantum Measurement
Let us then introduce three different pure states, the first one located in point $A$ (see Figure 4) represented by the vector $|\pi/3,\phi\rangle$, hence with polar angle $\theta=\pi/3$, the second one located in point $B$ (see Figure 4) represented by the vector $|2\pi/3,\phi\rangle$, hence with polar angle $\theta=2\pi/3$, and the third one located in the North Pole (see Figure 4), represented by the vector $|0,\phi\rangle$, (hence, this is the eigenstate describing Light, with a polar angle 00), assuming for simplicity that they all lie on a same plane, hence have the same asimuthal angle $\phi$. When a Light-Dark color-measurement is performed, the pre-measurement pure states deterministically transform into the fully decohered pre-collapse density states, obtained by plunging the associated point particles into the sphere, orthogonally with respect to the line subtended by the Light and Dark outcome states, which is the region of the percepts, i.e., of the ‘contextual reality’ relative to this specific color-measurement. This results for the pure states located in $A$, $B$ and $C$, respectively, in density states located in points $A^{\prime}$, $B^{\prime}$ and $C^{\prime}$, and given, respectively, by the density matrices
$\begin{pmatrix}1/4&0\\ 0&3/4\end{pmatrix},$ (42)
$\begin{pmatrix}3/4&0\\ 0&1/4\end{pmatrix}$ (43)
${\rm and}\quad\begin{pmatrix}1&0\\ 0&0\end{pmatrix}$ (44)
@@
If we consider the pure states $|{\pi\over 3},\phi\rangle$ and $|{2\pi\over 3},\phi\rangle$, located respectively in points $A$ and $B$, they are in the Light and Dark hemispheres of the Bloch sphere, hence they are two different colors, and we can now easily see that their transformation to the corresponding decohered density states, located respectively in points $A^{\prime}$ and $B^{\prime}$ (see Figure 4), and described respectively by the density matrices (42) and (43) exhibits a warping that is a dilation. Indeed, their distance is $1/3$, i.e., one third of the maximal distance between two stimuli, while the associated percepts, corresponding to the decohered quantum states, have a distance of $1/2$, i.e., one half of the maximal distance between two percepts. On the other hand, if we consider the pure states $|0,\phi\rangle$ and $|{\pi\over 3},\phi\rangle$, located respectively in points $C$ and $A$, belonging to the same color, namely the color Light, the opposite warping occurs. Indeed, on the stimuli side, the distance is again $1/3$, whereas on the percepts side, the corresponding density states being located in points $C^{\prime}$ and $A^{\prime}$, respectiveky, and described by the density matrices (44) and (42), we now have distance $1/4$. This means that a warping takes place which is now a contraction. This shows how the phenomenon of categorical perception is built into a the quantum measurement.
@@ 5 Visual Perception, Colors and Quanta
Let us analyze what the events are that correspond to this way of using the quantum measurement model to model the phenomenon of categorical perception of colors. And let us start from the situation where the stimulus, as measured by a physical measuring device that measures the frequency of electromagnetic radiation, is perceived by a human being, and from what we know takes place then.
@@
Suppose the light being looked at has a frequency of 595 Terahertz, given that 1 Terahertz equals $10^{12}$ Hertz, which corresponds to a wavelength of about 504 nanometers. If we consult one of the many color spectra of visible light available on the World-Wide Web, we can see that 504 nanometers wavelength corresponds to a color right in between green and blue, hence light of this frequency possesses a Blue-Green color, but it is not obvious whether when someone is asked to classify this light among one of the basic or prototype colors, whether that person will pick either Blue or Green. Each region that makes a transition from one basic color to another basic color contains a strip where it is really not clear which of the adjacent basic colors will be chosen if such a choice is imposed, for example in an experiment.
@@
That being said, we plan an experiment where we will show certain shades of color and on a Likert scale have participants choose to classify these shades of color with one of the basic colors. We will then investigate the extent to which we can model the shades of color in a complex Hilbert space as superpositions of the basis vectors each representing one of the basic colors. Let us note that in our example of the shade Blue-Green carried by light of frequency 595 Terahertz, this is a color carried by very pure light of exact frequency. We already noted that the prototype colors for the English language are ‘not’ the colors of the rainbow, for example the color Brown is not realized by light of a specific frequency. Colors of reflected light can represent very complex mixtures of different frequencies, just think of how white light is obtained by rotating a color disk rapidly, what happens there is very complex when the stimulus in question, the light shone from the rotating disk, is examined from physics. For what colors ‘are’, what happens in our eye on the retina, but especially in our brain, and perhaps just as importantly, in and with our mind, is fundamental. Given the eleven prototype colors considered as basic colors by Berlin & Kay (1969) and Rosch (1973), for our planned experiment we will also consider these colors as basic colors. This means that the set of all superpositions are the elements, more precisely the unit vectors, of an eleven-dimensional complex Hilbert space, with a basis of orthonormal vectors each of which represents one of the considered basic colors. Although we can also explicitly build the Bloch representation for this situation of eleven basic colors, and we plan to do so after we have collected the data from our experiment, we will not become explicit about it in this article, because the more than two-dimensional Bloch representations are more complicated, and we have focused, at least in part also for reasons of simplicity and intelligibility, but also because we wanted to turn to the fundamental in this article, to the situation of two colors, Light and Dark. After all, the fundamental, namely the way in which stimuli and expectations play a role in determining percepts, is equally and fully present in this situation of two colors.
@@
So let us return to the situation of two colors that we have called Light and Dark. Let it also be noted here that a restriction of vision to Light and Dark, and thus shades of gray, still leads to a powerful use of the sense of ‘sight’. Indeed, there are several animals that see only shades of gray, and this group is called the ‘monochromatic mammals’. Several bats, rodents and the common raccoon are among them. One common denominator among these mentioned species is that they are nocturnal animals, so their monochromatic vision gives them a special advantage at night. The shade of gray that contains the most uncertainty regarding its classability as Light or Dark is the shade that lies at the equator of the Bloch sphere (see Figure 4). It is plausible to associate a probability equal to 1/2 with the point at the center of the elastic that runs between North and South Poles in the extended Bloch model.
@@
Are the probabilities proposed by the Bloch model also obtained in the limit of the fractions obtained for the outcomes when an experiment is conducted to measure them? Regarding this question, we should mention the result of research published in Aerts & Sassoli de Bianchi (2015). If we imagine different individuals and the way they will decide to classify a particular shade of gray as Light rather than Dark or vice versa, it is very plausible to assume that there will be a fair amount of individual differences. Some will still classify grays that are above the equator as Dark while others will still classify grays that are below the equator as Light. It is shown in Aerts & Sassoli de Bianchi (2015) that the average over such individual differences tends to converge to a uniform distribution, which would then correspond to the quantum probabilities proposed in the previous sections of this paper. That is, the quantum measurement model would be a kind of first-order approximation to more complex measurement situations, which would then also explain why in cases where control over such individual differences is absent, it leads to good predictive models. It is beyond the scope of our current article to explain this complex situation of ‘quantum as first order approximation’ in more detail, we refer the interested reader to Aerts & Sassoli de Bianchi (2017) for a very detailed account of this situation.
Let us now come to our argument in connection with colors being quanta of light for human vision, and so for a raccoon these are only two colors Light and Dark, while for English-speaking humans they are eleven, White, Black, Red, Yellow, Green, Blue, Brown, Purple, Pink, Orange and Gray. Let us explain why these eleven colors are quanta of light for English-speaking humans. Suppose a friends has a bicycle colored with exactly the color of light corresponding to a frequency of 595 Terahertz. And suppose for a moment that a conversation develops about this bicycle, perhaps it is lost and they are looking for it. The color might then come up. Perhaps Green will be specified as the color when asked, or possibly Blue will be named. Or, when precision is important, Blue-Green is said. All three possibilities bring probative value to our claim that the eleven prototype colors are quanta of light for English-speaking people. Indeed, even if Blue-Green is given as the color for the bicycle to be found, this word Blue-Green is a combination of two other words, Blue and Green. Consider the pet-fish problem we mentioned in Section 4. Indeed, in quantum cognition, such combinations of words are described by superpositions of the vectors that describe the participles of these combinations (Aerts & Gabora, 2005a, b; Aerts et al., 2012). The cultural fact that it was decided in English to include these eleven designations as separate words in the English language, but not to introduce a separate word for the color that light of 595 Terahertz shines around, is what determines their status as quanta. Indeed, it is the dynamics produced by the phenomenon of categorical perception, the mechanism that causes certain stimuli, that is, in this case of colors, certain frequencies of light, to contract so that it becomes meaningful within the cognitive interaction to classify them all as Green, while other stimuli are pushed apart so that it becomes meaningful within the same cognitive interaction to classify them as different colors, Green and Blue, for example, that we wish to call ‘quantization’. It is our hypothesis that the same contraction and pushing apart occurs in the quanta that light shows to physical measuring devices and their cognitive activities, so that, in the case of light, the photons emerge as quanta. The process of quantization ends, or better stabilizes, as an optimum of this cognitive interaction.
This means that as far as human cognition is concerned, the interaction that takes place between the stimulus, on the one hand, and the expectation pattern of the person perceiving, on the other hand, in order to arrive at perception, is of a greater complexity than the experiment with the tricked cards described in brunerpostman1949 suggests. It is likewise of greater complexity than the dynamics of elastics breaking as we suggested in the extended Bloch model. That does not mean, however, that these models are irrelevant to making progress in the depth of what human perception and cognition are. Indeed, they both contain essential elements of this human perception and cognition and suggest that more complex interactions need to be expected. But, the result demonstrated in Aerts & Sassoli de Bianchi (2015, 2017), makes it clear that such a more complex situation of interaction between the measuring device and the entity being measured should not necessarily lead to the need for a more complex model of measurement, at least not if we are satisfied with the linear model proposed by quantum mechanics being a first order approximation of what takes place during measurement. So it is especially in comparison to a classical description that assumes that no interaction at all takes place between measuring device and the entity being measured that the first order approximation of the linear quantum model is an essential step forward.

### v2

@@ 3 Quantum Measurement and Categorical Perception
We introduce three pure states. The first one located in point $A$ (see Figure 2) with polar angle $\theta=\pi/3$, the second one located in point $B$ (see Figure 2) with polar angle $\theta=2\pi/3$ and the third one located in the north pole (see Figure 2), (hence, this is the eigenstate describing Light, with polar angle $\theta=0$), assuming for simplicity that they all lie on the same plane, hence have the same asimuthal angle $\phi$. When a Light-Dark color-measurement is performed, the pure states deterministically transform into the fully decohered density states, obtained by plunging the associated points into the sphere, orthogonally with respect to the line subtended by the Light and Dark outcome states, which is the region of the percepts relative to this specific color measurement. This results for the pure states located in $A$, $B$ and $C$, respectively, in density states located in the points $A^{\prime}$, $B^{\prime}$ and $C^{\prime}$.
@@
If we consider the pure states located respectively in points $A$ and $B$, they are in the Light and Dark hemispheres of the Bloch sphere, hence they are two different colors, and we can now easily see that their transformation to the corresponding decohered density states, located respectively in points $A^{\prime}$ and $B^{\prime}$ (see Figure 2) and described by density matrices exhibits a warping that is a dilation. In fact, their distance is $1/3$, that is, one third of the maximal distance between two stimuli, while the associated percepts, corresponding to the decohered quantum states, have a distance of $1/2$, that is, one half of the maximal distance between two percepts.
On the other hand, if we consider the pure states located respectively in points $C$ and $A$, belonging to the same color, namely the color Light, the opposite warping occurs. Indeed, on the stimuli side, the distance is again $1/3$, whereas on the percepts side, the corresponding density states being located in points $C^{\prime}$ and $A^{\prime}$, respectiveky, and described by density matrices, we now have distance $1/4$. This means that a warping takes place, which is now a contraction.
In Figure 3 we look in more detail at the effects of contraction or dilation, and have divided the $\pi$ radians or $180$ degrees circle arc connecting the North Pole with the South Pole into $18$ equal parts, each part spanning an angle of $0.05556\cdot\pi$ radians or $10$ degrees. We thus determine 18 points on the great circle connecting the North Pole and the South Pole where each two consecutive points are $0.05556\cdot\pi$ radians or $10$ degrees apart. In a region localized around the North Pole and a similar region localized around the South Pole, the decoherence of a quantum measurement causes the pure states associated with each of these points to be transformed into density states that are closer to each other, and thus a contraction occurs in these regions. In the table of Figure 3, these cases are indicated in green. In a region located between the North pole and the South pole, centered around the point $\frac{1}{2}\pi$ radians away from both, the decoherence of quantum measurement causes dilation. The pure states connecting to the points there are transformed into density states farther apart. In the table of Figure 3, these cases are indicated in blue.
Figure 3: A more detailed representation of the contraction or dilation that takes place in a quantum measurement. We have divided the $pi$ radians long circle arc connecting the North Pole with the South Pole in equal parts, each part spanning an angle of $0.05556\cdot\pi$ radians or $10$ degrees. We thus determine $18$ points on the great circle arc connecting the North Pole and South Pole. In a region localized around the North Pole and a similar region localized around the South Pole, the decoherence of a quantum measurement causes the pure states associated with each of these points to be transformed into density states that are closer to each other and thus a contraction occurs. In the table, these cases are indicated in green. In a region located between the North Pole and the South Pole centered around the point $\frac{1}{2}\cdot\pi$ radians away from both, the decoherence of the quantum measurement causes a dilation. The pure states connecting to the points there are transformed into density states farther apart. In the table, these cases are indicated in blue.
This shows how the phenomenon of categorical perception is built into the quantum measurement.
@@ 4 Visual Perception, Colors and Quanta
Let us analyze the events that correspond to this way of using the quantum measurement model to model the phenomenon of categorical perception of colors. And let us start from the situation where the stimulus, as measured by a physical measuring device that measures the frequency of electromagnetic radiation, is perceived by a human being, and from what we know takes place then.
Suppose that the light being looked at has a frequency of 595 terahertz, given that 1 terahertz equals $10^{12}$ hertz, which corresponds to a wavelength of about 504 nanometers.
Figure 4: A representation of the colors of visible light corresponding to the wavelengths of that light. Light with wavelength 504 nanometer has a color right between green and blue.
@@
If we consult one of the many color spectra of visible light available on the World Wide Web, we can see that the 504 nanometer wavelength corresponds to a color right between green and blue (see Figure 4), so light of this frequency possesses a Blue-Green color, but it is not obvious whether when someone is asked to classify this light among one of the basic or prototype colors, whether that person will pick either Blue or Green. Each region that makes a transition from one basic color to another basic color contains a strip in which it is really not clear which of the adjacent basic colors will be chosen if such a choice is imposed, for example, in an experiment.
@@
That being said, we plan an experiment where we will show certain shades of color and on a Likert scale have participants choose to classify these shades of color with one of the basic colors. We will then investigate the extent to which we can model the shades of color in a complex Hilbert space as superpositions of the basis vectors, each representing one of the basic colors. Let us note that in our example of the shade Blue-Green carried by light of frequency 595 terahertz, this is a color carried by very pure light of exact frequency. We already noted that the prototype colors for the English language are ‘not’ the colors of the rainbow, for example, the color Brown is not realized by light of a specific frequency. Colors of reflected light can represent very complex mixtures of different frequencies, just think of how white light is obtained by rotating a color disk rapidly. What happens there is very complex when the stimulus in question, the light shone from the rotating disk, is examined from physics. With respect to the nature of colors, what happens in our eye on the retina, but especially in our brain, and perhaps just as importantly, in and with our mind, is fundamental. Given the eleven prototype colors considered basic colors by Berlin & Kay (1969) and Rosch (1973), for our planned experiment we will also consider these colors as basic colors. This means that the set of all superpositions is the set of elements, more precisely, the unit vectors, of an eleven-dimensional complex Hilbert space, with a basis of orthonormal vectors, each of which represents one of the considered basic colors. Although we can also explicitly build the Bloch representation for this situation of eleven basic colors, and we plan to do so after we have collected the data from our experiment, we will not become explicit about it in this article because the more than two-dimensional Bloch representations are more complicated, and we have focused, at least in part also for reasons of simplicity and intelligibility, but also because we wanted to turn to the fundamental in this article, to the situation of two colors, Light and Dark. After all, the fundamental, namely, the way in which stimuli, expectations, and measurements play a role in determining percepts, is equally and fully present in this situation of two colors.
@@
So, let us return to the situation of two colors that we have called Light and Dark. Let it also be noted here that a restriction of vision to Light and Dark, and thus shades of gray, still leads to a powerful use of the sense of sight. In fact, there are several animals that see only shades of gray, and this group is called the ‘monochromatic mammals’. Several bats, rodents, and the common raccoon are among them. One common denominator among these mentioned species is that they are nocturnal animals, so their monochromatic vision gives them a special advantage at night. The gray shade that contains the most uncertainty regarding being classified as Light or Dark is the shade that lies at the equator of the Bloch sphere (see Figure 2). It is plausible to associate a probability equal to 1/2 with the point at the center of the line that runs between the north and south poles in the extended Bloch model.
@@
Are the probabilities proposed by the Bloch model also obtained in the limit of the fractions obtained for the outcomes when an experiment is conducted to measure them? Regarding this question, we should mention the result of research published in Aerts & Sassoli de Bianchi (2015). If we imagine different individuals and the way they will decide to classify a particular shade of gray as Light rather than Dark or vice versa, it is very plausible to assume that there will be a fair amount of individual differences. Some will still classify grays that are above the equator as Dark while others will still classify grays that are below the equator as Light. It is shown in Aerts & Sassoli de Bianchi (2015) that the average over such individual differences tends to converge to a uniform distribution, which would then correspond to the quantum probabilities. This means that the quantum measurement model would be a kind of first-order approximation to more complex measurement situations, which would then also explain why in cases where control over such individual differences is absent, it leads to a good predictive model. It is beyond the scope of our current article to explain this complex situation of quantum as first-order approximation in more detail, and we refer the interested reader to Aerts & Sassoli de Bianchi (2017) for a very detailed account of this situation.
Let us now come to our argument in connection with colors being quanta of light for human vision, and so for a raccoon these are only two colors Light and Dark, while for English-speaking humans they are eleven, White, Black, Red, Yellow, Green, Blue, Brown, Purple, Pink, Orange and Gray. Let us explain why these eleven colors are quanta of light for English-speaking humans. Consider the pet-fish problem we mentioned in Section 2. In fact, in quantum cognition, such combinations of words are described by superpositions of the vectors that describe the participles of these combinations (Aerts & Gabora, 2005a, b; Aerts et al., 2012).
If we claim that our analysis makes it possible to consider colors as quanta for human vision, we must emphasize that there is an essential aspect of ‘human language’ present. In fact, also in the anthropological research of Berlin & Kay (1969) the focus is on the existence of words that denote colors in a given language. That aspect was also already crucial in the situation of the Dani, as studied by Rosch Heider (1971, 1972); Rosch (1973), they possessed only two words denoting colors. In this sense, our observation that colors are quanta for the human vision of light is a special case of what is claimed in Aerts & Beltran (2022a), that ‘words are the quanta of human language’. Is it really the ‘words’, which are the quanta of human language? Well, no, it is the underlying ‘concepts’, as we made clear in Aerts & Beltran (2022a), and likewise in Rosch Heider (1971, 1972); Rosch (1973), it was all about the underlying concepts through which, incidentally, her research led to the introduction of the prototype theory for concepts. Is there no quantization before the emergence of language? This is an interesting question on which we want to dwell for a moment. There is no scientific consensus on the origin of human language, but one of the most plausible views postulates that the first languages consisted of gestures (Tomasello,, 1996; Pika & Mitani,, 2006). These gestures, with the disadvantage of making it difficult to communicate with someone in a more intimate way, evolved into movements accompanied by sounds more inward in the body, the mouth, lips, tongue, and larynx. Words and what we now commonly understand by language came much later, when more complex abstractions are communicated. It is our opinion that the phenomenon of categorical perception played a role from the very beginning, and probably thus then at the level of gestures, in instilling conceptual accumulations leading to clumping and finally quanta as a stabilization of this part of the process as a consequence of the contractions present on the stimuli and distinctions between the quanta as a consequence of the dilations.
The time has come to clarify another aspect of our analysis. It rests on the consideration of the quantum measurement process that makes us interact with physical reality by means of measuring devices, on the one hand, and the consideration of human perception, which makes us interact with human conceptual reality by means of the human perceptual device, on the other hand, and then to consider both as special cases of a more general situation where measurement takes place on entities of a reality. In philosophy of mind, the notion of ‘qualia’ was introduced to denote the subjective sensation of a human experience (Lewis,, 1929). Although there is no general consensus on this notion, neither among philosophers nor among scientists who study the human mind, such as neurologists, with some even doubting its existence, here we wish to use the notion in a specific way. In doing so, we do not pretend to resolve any controversy surrounding it, our intention is only to show that the notion of qualia can be used to shed additional light on the analysis we propose in this article about quantum measurement versus human perception. In this sense we claim that in a similar way that we can argue that ‘colors are quanta of light for human vision’, we can argue that ‘photons are qualia for physical measuring devices that can identify them’.